# long-var-wrap = dv

### v2

@@ IV.1 Discrete Symmetry
where $\ell=\sqrt{2/|r|}$ is the domain wall thickness, which results from a trade-off between the surface tension term $\nabla^{4}\varphi$ and the potential term $r\nabla^{2}\varphi$. This domain wall profile is, in fact, valid for any average composition $\bar{\varphi}$. As discussed in the previous chapter, there may be more than two phases, but the walls between them will still have a hyperbolic tangent profile with the same length scale $\ell$. See Ref. [59] for a more complete discussion.
@@
Figure 7: The domain wall of thickness $\ell$ with the profile of the hyperbolic tangent, light blue, compared to numerical solutions for small $\alpha$ and $\bar{\varphi}$. Parameters are $u=-r=10$, $\alpha=2$, $(\bar{\varphi}_{1},\bar{\varphi}_{2})=-(0.1,0.2)$.
For some restricted versions of the NRCH, one may construct a “spurious gradient flow” [60, 61]. These methods, however, are not applicable in our case, due to the symmetries chosen. They balance “spurious” pressure and chemical potential across a stationary domain wall; however, this is not available for traveling waves or domain walls. A model similar to those considered in  [60, 61] is obtained by adding a term $-\rho\varphi_{1}\varphi_{2}$ to the $\mathrm{C}_{4}$ model, breaking its symmetry down to $\mathbb{Z}_{2}$. As seen in recent investigations [62], in models with only $\mathbb{Z}_{2}$-symmetry the onset of traveling-wave states happens for finite non-reciprocity [63, 62]. This thus illustrates how symmetries allow for achieving much “cheaper” non-equilibrium steady states; however, an arbitrarily small $\alpha$ still results in traveling states. Similar behavior was observed in Ref. [59].
@@ IV.2 Continuous Symmetry
We can find an exact solution for the $\mathrm{SO}(2)$ model in the case of $\bar{\varphi}=0$, corresponding to a slice of the linear stability diagram of the last section where the entire plane, for $r<0$, is unstable [54, 51]. This solution has the form
$\varphi_{1}(t,\bm{x};\bm{k})$ $=A(k)\sin\phi(t,\bm{x};\bm{k}),$ (19a)
$\varphi_{2}(t,\bm{x};\bm{k})$ $=A(k)\cos\phi(t,\bm{x};\bm{k}),$ (19b)
$\phi(t,\bm{x};\bm{k})$ $=\bm{k}\cdot\bm{x}-\omega(k)t.$ (19c)
@@
$\omega(k)=\alpha k^{2}\Gamma,\quad uA(k)^{2}=-r-k^{2}.$ (20)
@@
Note that $\varphi_{a}(t,\bm{x};\bm{k})$ solves the full non-linear equations of motion and is only possible due to the rotational symmetry of these equations. The solution is illustrated in Figure 8. The plot on the left shows the fields in real space, in the direction parallel to the wave vector $\bm{k}$. We may also visualize the configuration in $(\varphi_{1},\varphi_{2})$-space by following the line parametrized by $x$, shown on the right, where the minimum of the potential is indicated by dashed black lines. This can be thought of as mapping the circle $x\in S^{2}$ into field space, with shorter wavelength solutions corresponding to the mapping winding around the origin more times.
Figure 8: The traveling wave solution for $\bar{\varphi}=0$. To the left, the fields are plotted in one dimension of real space, parallel to the wave vector. On the right, the same configuration is shown in field space.
Linear stability analysis from substituting $\varphi_{a}(t,\bm{x})=\varphi_{a}(t,\bm{x};\bm{k})+\delta\varphi_{a}(\bm{x})$ into Eq. (5), as detailed in Appendix A, reveals that although there may be unstable higher $\bm{k}$ solutions, the longest wavelength solution is always stable. The velocity of the wave is controlled by the non-reciprocal parameter, with $\alpha=0$ corresponding to continuous phase separation. As $\alpha$ increases, the wave starts to move. This traveling solution is an example of a non-equilibrium steady-state solution [51].
Decreasing further below $r=0$ yields shorter wavelength solutions. Minimizing the free energy density $\mathcal{F}=-\frac{3}{4}uA^{2}$ corresponds to maximizing the amplitude of the wave. The global minimum is the state with the longest wavelength $k=2\pi/L$ where $L$ is the system size, while solutions with higher wave numbers are metastable states, at least for $\alpha=0$. Assuming $|r|\gg k^{2}$, the amplitude of the wave is $A\approx\sqrt{-r/u}$, corresponding to mapping the circle in field space around the bottom of the potential, as illustrated in Figure 8. In the metastable states for which $k=2\pi n/L$ and $n>1$ the circle is mapped $n$ times around the brim of the hat. As $n$ increases, the amplitude of the wave decreases as the kinetic term of the free energy acts as a tension in the line. This solution is closely related to the topological defects this model admits [52], as far away from the core of a defect, the system forms a traveling sine-wave. The general analysis holds for small finite $\bar{\varphi}$, see video_2.mp4 in[56] for further illustration. When $|r|<k^{2}$, Eq. (19) no longer corresponds to a stationary solution of the equations of motion, neither stable, nor metastable, nor unstable, as the tension has become too high.
@@
Figure 9: The approximate solution for $\bar{\varphi}\approx\varphi^{\prime}$ is shown top left in gray, while the numerical solution is overlayed in red and black. Bottom left shows the deviation between analytical and numerical solutions, including noise $D=2\times 10^{-4}$. The right panel shows the solution in $\varphi_{i}$ space. Parameters are $u=-r=40$, $(\varphi_{1},\varphi_{2})=-(0.8,0)$.
@@
$\varphi_{1}(t,\bm{x};\bm{k})$ $=\delta\cos 2\phi(\bm{x};\bm{k})-\bar{\varphi},$ (21a)
$\varphi_{2}(t,\bm{x};\bm{k})$ $=2\sqrt{\bar{\varphi}\delta}\cos\phi(\bm{x};\bm{k}),$ (21b)
$\phi=\bm{k}\cdot\bm{x},\quad\varphi^{\prime}$ $=\sqrt{|r|/u},\quad\delta=\varphi^{\prime}-\bar{\varphi}.$ (21c)
We show in Appendix B that for $\delta\ll\varphi^{\prime}$ and in the long wavelength limit, this is a good approximation of the stationary solution. While the previous solution corresponds to mapping a line parallel to $\bm{k}$ around the brim of the hat, this solution can be visualized as mapping this same loop to a small sector of the brim. This is illustrated in Figure 9, where the analytical solution is plotted in gray, together with the numerical solution and the deviation between them. The relaxation of the system towards this solution is illustrated in video_3.mp4[56].
Videos 2 and 3 are simulations of the NRCH model with an added noise term od the form $\sqrt{2D}\,\xi_{a}(\bm{x},t)$, where $\xi$ represents Gaussian white noise of unit strength with zero mean. The noise-strength in the simulations is set to $D=2\times 10^{-4}$ in the appropriate non-dimensionalized form. The value is chosen to show the robustness of the observed behavior to weak noise, while still illustrating the deviation away from the analytical solution of Figure 9.
@@ IV.3 Relation to the Complex Ginzburg-Landau equation
[…]
has similarities to the Complex Ginzburg-Landau equation (CGLE) [29]. One may consider CGLE as the non-reciprocal extension of model A [65], and the $\mathrm{SO}(2)$ NRCH, Eq. (5), as the non-reciprocal extension of model B. The CGLE may include higher-order non reciprocal terms, which have been explored recently for the NRCH [54]. In the case of the CGLE, these term are necessary in order to obtain non-trivial behavior, as the liner term can be removed by a time-dependent field transformation, $\varphi_{a}\rightarrow[e^{-\alpha t\epsilon}]_{ab}\varphi_{b}$. At a formal level, one might, naively, extend this argument to the $\mathrm{SO(2)}$ NRCH as well, with the transformation $\varphi_{a}\rightarrow[e^{\alpha\nabla^{2}t\epsilon}]_{ab}\varphi_{b}$. Applying this transformation to the traveling-wave solution Eq. (19) gives a static-wave solution that solves the $\alpha=0$ equation of motion. This would hint at a hidden $\mathbb{Z}_{2}$ symmetry, so that the full symmetry is $\mathrm{O}(2)\simeq\mathrm{SO}(2)\rtimes\mathbb{Z}_{2}.$
We note, however, that the conservation law introduces intricacies into the structure of the equation that hampers the applicability of this argument, as the Laplacian operator in the exponential renders the transformation highly non-local. For example, the NRCH admits a broken-symmetry state in two dimensions [51], which is not the case case for model B. Furthermore, while Eq. (22) admits a Gaussian steady state, as one would expect since it can be mapped to an equilibrium theory, the same is not the case for Eq. (5). This indicates that the transformation is in general not well-defined.
@@ V Conclusion
In this article, we have explored the phase behavior of the non-reciprocal Cahn-Hilliard model. By choosing two reduced models—one with rotational invariance in field space and one without—we can fully characterize the linear behavior and see the effect of continuous symmetry on phase separation. The main behaviors observed are homogeneous configurations, static phase separation—continuous or with a well-defined domain wall—and non-equilibrium steady states with non-zero currents. Using this construction we have shown the effect of a continuous and connected symmetry on phase separation, where there is no well-defined domain wall. This effect remains in the case of traveling wave states where the phases chase each other due to non-reciprocal interactions.
@@ Acknowledgements.
The Authors would like to thank Giulia Pisegna and Jacopo Romano for stimulating discussions. We acknowledge support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.

### v3

@@ IV.1 Discrete Symmetry
where $\ell=\sqrt{2/|r|}$ is the domain wall thickness, which results from a trade-off between the surface tension term $\nabla^{4}\varphi$ and the potential term $r\nabla^{2}\varphi$. This domain wall profile is, in fact, valid for any average composition $\bar{\varphi}$. Since there are four degenerate minima in the potential landscape, as illustrated in Fig. 1, there may be more than two phases, but the walls between them will still have a hyperbolic tangent profile with the same length scale $\ell$. See Ref. [50] for a more complete discussion.
@@
Figure 6: Numerical simulation of traveling phase-separated domains in the model with discrete symmetry, for small $\alpha$ and $\bar{\varphi}$, at two different times $T$. The concentrations of the two different species are illustrated in full red and black lines, the dashed lines illustrate the corresponding value for $\bar{\varphi}$. The domain wall of thickness $\ell$, with a hyperbolic tangent profile, is illustrated in light blue. The parameters are u=r=10u=-r=10, $\alpha=2$, $(\bar{\varphi}_{1},\bar{\varphi}_{2})=-(0.1,0.2)$.
For some restricted versions of the NRCH, one may construct a “spurious gradient flow” [61, 60]. They balance “spurious” pressure and chemical potential across a stationary domain wall; however, this is not available for traveling waves or domain walls. A model similar to those considered in  [61, 60] is obtained by adding a term $\rho\varphi_{1}\varphi_{2}$ to the $\mathrm{C}_{4}$ model, breaking its symmetry down to $\mathbb{Z}_{2}$. As seen in recent investigations [62], in models with only $\mathbb{Z}_{2}$-symmetry the onset of traveling-wave states happens for finite non-reciprocity [43, 63, 62]—when the non-reciprocal coupling $\alpha$ is larger than the reciprocal coupling $\rho$ and the system passes an exceptional point. This thus illustrates how symmetries allow us to achieve steady-states with a non-zero current without the need to turn up the non-reciprocal coupling beyond a finite threshold. We can relate this behavior to the symmetry of the model [^3] Here, the special point corresponds to the reciprocal coupling $\rho$ vanishing. This corresponds to a symmetry $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ under $\varphi_{a}\rightarrow-\varphi_{a}$, $\alpha\rightarrow-\alpha$..
[…]
This observation opens up further avenues for investigating the effects of symmetry on the NRCH model. The two models could be linked by adding a term $g\varphi_{1}^{2}\varphi_{2}^{2}$, to investigate the transition between the different regimes, and the effects of small symmetry breaking. Furthermore, the $C_{4}$ model can be extended to $C_{n}$, and one would expect the to models to converge as $n\rightarrow\infty$. Another generalization would consider symmetry in models with non-reciprocal interactions between larger number of fields, such as in the models studied in Ref. [68].
@@ IV.2 Continuous Symmetry
We can find an exact solution for the $\mathrm{SO}(2)$ model in the case of $\bar{\varphi}=0$, corresponding to a slice of the linear stability diagram of the last section where the entire plane, for arbitrary $r<0$, is unstable [53, 51]. We obtain this solution by making a plane-wave ansatz as follows:
$\varphi_{1}(\bm{x},t;\bm{k})$ $=A(k)\sin\phi(\bm{x},t;\bm{k}),$ (18a)
$\varphi_{2}(\bm{x},t;\bm{k})$ $=A(k)\cos\phi(\bm{x},t;\bm{k}),$ (18b)
$\phi(\bm{x},t;\bm{k})$ $=\bm{k}\cdot\bm{x}-\omega(k)t.$ (18c)
@@
$\omega(k)$ $=\alpha k^{2},$ $uA(k)^{2}$ $=-r-k^{2}.$ (19)
@@
The function $\phi$ represents the phase of the wave, which depend on $\bm{x}$ and $t$. The solution corresponds to the initial phase difference of the two species of $\pi/2$. If the initial condition for the phase difference is not set to $\pi/2$, it will relax to $\pi/2$ with a relaxation time that is controlled by the wavevector of the solution, as shown in Ref. [51]. Note that $\varphi_{a}(t,\bm{x};\bm{k})$ solves the full non-linear equations of motion and is only possible due to the rotational symmetry of these equations. The solution is illustrated in Figure 7(a). The plot on the left shows the fields in real space, in the direction parallel to the wave vector $\bm{k}$. We may also visualize the configuration in $(\varphi_{1},\varphi_{2})$-space by following the line parametrized by $x$, shown on the right, where the minimum of the potential is indicated by dashed black lines. This can be thought of as mapping the circle $x\in S^{1}$ into field space, with shorter wavelength solutions corresponding to the mapping winding around the origin more times. From the discussion in the previous section, we may explicitly evaluate the entropy production for the solution given in Eq. (18):
$\dot{\sigma}=\frac{\alpha^{2}}{uD}k^{2}(|r|-k^{2}).$ (20)
@@
Figure 7: (a) The traveling wave solution for $\bar{\varphi}=0$. To the left, the fields are plotted in one dimension of real space, parallel to the wave vector. On the right, the same configuration is shown in field space. (b) The approximate solution for $\bar{\varphi}\approx\varphi^{\prime}$ is shown top left in gray, while the numerical solution is overlayed in red and black. Bottom left shows the deviation between analytical and numerical solutions, including noise $D=2\times 10^{-4}$. The right panel shows the solution in $\varphi_{i}$ space. Parameters are u=r=40u=-r=40, $(\bar{\varphi}_{1},\bar{\varphi}_{2})=-(0.8,0)$.
To establish the stability of this solution, we write the solution in a polar form and consider perturbations in the radial and tangential directions. This is detailed in Appendix A. After changing to a rotating frame of reference, we obtain a time- and space-independent system of equations, which can be used for a stability analysis. We find that to ensure stability the criterion $k^{2}<\frac{1}{3}|r|$ needs to hold. Therefore, for a sufficiently large system, the longest wavelength solution is always stable. The velocity of the wave is controlled by the non-reciprocal parameter, with $\alpha=0$ corresponding to continuous phase separation. As $\alpha$ increases, the wave starts to move. This traveling solution is an example of a non-equilibrium steady-state solution [51].
Decreasing further below $r=0$ yields shorter wavelength solutions. Minimizing the free energy density $\mathcal{F}=-\frac{3}{4}uA^{2}$ corresponds to maximizing the amplitude of the wave. The global minimum is the state with the longest wavelength $k=2\pi/L$ where $L$ is the system size, while solutions with higher wave numbers are metastable states, at least for $\alpha=0$. Assuming $|r|\gg k^{2}$, the amplitude of the wave is $A\approx\sqrt{-r/u}$, corresponding to mapping the circle in field space around the bottom of the potential, as illustrated in Figure 7(a). In the metastable states for which $k=2\pi n/L$ and $n>1$ the circle is mapped $n$ times around the brim of the hat. As $n$ increases, the amplitude of the wave decreases as the kinetic term of the free energy acts as a tension in the line. This solution is closely related to the topological defects this model admits [52], as far away from the core of a defect, the system forms a traveling sine-wave. The general analysis holds for small finite $\bar{\varphi}$, see video_2.mp4 in[57] for further illustration. When $|r|<k^{2}$, Eq. (18) no longer corresponds to a stationary solution of the equations of motion, neither stable, metastable, nor unstable, as the tension has become too high.
@@
$\varphi_{1}(\bm{x},t;\bm{k})$ $=\delta\cos 2\phi(\bm{x};\bm{k})-\bar{\varphi},$ (21a)
$\varphi_{2}(\bm{x},t;\bm{k})$ $=2\sqrt{\bar{\varphi}\delta}\cos\phi(\bm{x};\bm{k}),$ (21b)
$\phi=\bm{k}\cdot\bm{x},\hskip 42.67912pt\varphi^{\prime}$ $=\sqrt{|r|/u},\hskip 42.67912pt\delta=\varphi^{\prime}-\bar{\varphi}.$ (21c)
We show in Appendix B that for $\delta\ll\varphi^{\prime}$ and in the long wavelength limit, this is a good approximation of the stationary solution. While the previous solution corresponds to mapping a line parallel to $\bm{k}$ around the brim of the hat, this solution can be visualized as mapping this same loop to a small sector of the brim. This is illustrated in Figure 7(b), where the analytical solution is plotted in gray, together with the numerical solution and the deviation between them. The relaxation of the system towards this solution is illustrated in video_3.mp4 [57].
Videos 2 and 3 are simulations of the NRCH model with an added noise term of the form $\sqrt{2D}\,\bm{\nabla}\cdot\bm{\xi}_{a}(\bm{x},t)$, where $\bm{\xi}_{a}$ represents Gaussian white noise of unit strength with zero mean. The noise-strength in the simulations is set to $D=2\times 10^{-4}$ in the appropriate non-dimensionalized form. The value is chosen to show the robustness of the observed behavior to weak noise, while still illustrating the deviation away from the analytical solution of Figure 7(b). The solutions presented correspond to infinite systems or systems with periodic boundary conditions. We expect the bulk properties to hold for other boundary conditions for sufficiently large system sizes. The effect of confinement can be implemented by incorporating a lower bound on the values of the wavevector, as a first approximation.
Video 2 illustrates that a traveling wave state close to that of Eq. (18) exists even for $\bar{\varphi}\neq 0$. The video also illustrates the possibility of topological defects in this system, which may occur due to the initial conditions or added noise. In one dimension, these appear as point defects, with traveling-wave domains between them. When the system is initiated at random initial conditions and the coupling is relatively weak, topological defects appear and form an arrested structure that disrupts global order, while sufficiently strong non-reciprocal coupling suppresses topological defects and creates global (polar) order [52], which is robust with respect to noise, even in two dimensions [51]. While weak noise is not expected to fundamentally change the picture provided by the linear stability analysis as long as defects are suppressed, strong noise can lead to nontrivial qualitative changes in the behavior of the system. A further possible complication is the appearance of pseudo-resonances [70]. Due to the existence of non-reciprocal interaction, the linear-stability matrix can be non-normal for certain parameter-values. This can have the effect that even for systems with eigenvalues where the real part is negative (such that all linear perturbations ultimately decay) some modes (in the pseudo-spectrum) feature transient growth, thereby significantly amplifying perturbations, albeit remaining finite in the long time limit [71]. We note that nonlinear non-reciprocal terms have been shown to lead to the emergence of effervescence, as well as self-generated noise due to spatiotemporal chaos [53].
@@ IV.3 Relation to the Complex Ginzburg-Landau equation
$\partial_{t}\varphi_{a}=-\left[{\frac{\partial\mskip 0.0mu\mathcal{V}_{\mathrm{cs}}}{\partial\varphi_{a}}}-\nabla^{2}\varphi_{a}+\alpha\epsilon_{ab}\varphi_{b}\right],$ (22)
has similarities to the Complex Ginzburg-Landau equation (CGLE) [29]. One may consider CGLE as the non-reciprocal extension of model A [72], and the $\mathrm{SO}(2)$ NRCH, Eq. (5), as the non-reciprocal extension of model B. The CGLE may include higher-order non reciprocal terms, which have been explored recently for the NRCH [53]. In the case of the CGLE, these terms are necessary in order to obtain non-trivial behavior, as the linear term can be removed by a time-dependent field transformation, $\varphi_{a}\rightarrow[e^{-\alpha t\epsilon}]_{ab}\varphi_{b}$. One might, naively, extend this argument to the $\mathrm{SO(2)}$ NRCH as well, with the transformation $\varphi_{a}\rightarrow[e^{\alpha\nabla^{2}t\epsilon}]_{ab}\varphi_{b}$. Applying this transformation to the traveling-wave solution Eq. (18) gives a static-wave solution that solves the $\alpha=0$ equation of motion. This would hint at a hidden $\mathbb{Z}_{2}$ symmetry, so that the full symmetry is $\mathrm{O}(2)\simeq\mathrm{SO}(2)\rtimes\mathbb{Z}_{2}.$ However, the $\varphi^{2}$ term is no longer invariant of the transformation when including the Laplacian operator in the exponential, so the transformation cannot be used to eliminate the linear non-reciprocal term. Indeed, such a transformation would render the NRCH with only a linear non-reciprocal term more-or-less equivalent to model B. However, this model admits a broken-symmetry state in two dimensions [51], which is not the case for model B. Furthermore, while Eq. (22) admits a Boltzmann distribution as a steady state, as one would expect since it can be mapped to an equilibrium theory, the same is not the case for Eq. (5).
A similar observation was made in Ref. [73]—the conserved version of the KPZ was at first believed to obey a modified version of the Galilean symmetry. This symmetry is protects the KPZ non-linearity from renormalization, and is thus crucial for the existence of the KPZ universality class. However, more careful analysis shows that this is not true in the conserved case, and the non-linearity is indeed renormalized [73], illustrating how careful analysis of symmetries are crucial in non-equilibrium models when considering conservation laws.
@@ V Conclusion
In this article, we have explored the phase behavior of the non-reciprocal Cahn-Hilliard model. By choosing two reduced models—one with rotational invariance in field space and one without—we can fully characterize the linear behavior and see the effect of continuous symmetry on phase separation. The main behaviors observed are homogeneous configurations, static phase separation—continuous or with a well-defined domain wall—and non-equilibrium steady states with non-zero currents. Using this construction we have shown the effect of a continuous symmetry on phase separation, where there is no well-defined domain wall. This effect remains in the case of traveling wave states where the phases chase each other due to non-reciprocal interactions.
@@ Acknowledgements.
The Authors would like to thank Giulia Pisegna, Jacopo Romano, and Suropriya Saha for stimulating discussions. We acknowledge support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.